OpenStudy (anonymous):

Each year, Terrell buys a plane ticket to visit his grandmother for a week. Last year, he paid one hundred thirty-five dollars. This year, he paid one hundred eighty-nine dollars. What was the percent increase in the cost of his plane ticket? Show your work.

OpenStudy (anonymous):

increased amount = 189-135 =54 % inreased = 54/135*100 = 40%